(_________,1/4) is a solution of y= 3x-7
Answer by mananth(16949)

(__2/3_______,1/4) is a solution of y= 3x-7
..
plug the value 1/4 for y
1/4
= 3x-7
multiply by 4
1=4(3x-7)
1=12x-7
+7 +7
8=12x
/12
8/12 = x
x= 2/3
...